Compute the direction and magnitude of the resultant of the following system of coplanar and concurrent forces; 20 N, 40 N, 25N, 42N, and 12 N making an angle of 30
1
Expert's answer
2021-08-09T08:04:08-0400


In unit vector form

"F_x = 20 cos 30 - 40 cos 60 -25 +12 cos 45 = - 19.194 N\\
F_y = 20 sin 30 - 40 sin 60 -42 +12 sin 45 = - 5.844 N\\
F= \sqrt{F_x^2+F_y^2}\\
F= \sqrt{(-19.194)^2+(-5.844)^2}\\
F= 20.064 N\\
\alpha = \tan^{-1}(\frac{-5.844}{-19.194})\\
\alpha = 16.934^0\\"

Therefore, the angle is 1800+16.9340 = 196.9340